From: Jet Propulsion Laboratory
Posted: Wednesday, July 18, 2001
Wednesday, July 18, 2001 (DOY 192/19:00:00 to DOY 199/19:00:00 UTC)
Launch / Days since Launch = Nov. 7, 1996 / 1715 days
Start of Mapping / Days since Start of Mapping = April 1, 1999 / 839 days
Total Mapping Orbits = 10,556
Total Orbits = 12,239
Background Sequences - The spacecraft is operating nominally in performing daily recording and transmission of science data. The mm150 sequence has performed well since it started on 01-172 (6/21/01). It terminates on 01-199 (7/18/01). The mm151 sequence, successfully uplinked on 01-197 (7/16/01), begins executing on 01-200 (7/19/01).
Other - Dust storms spreading across the planet are making narrow angle photographic imaging of Mars impracticable. Therefore, ROTO operations remain suspended until the global dust storms subside. MGS has completed 155 ROTOs to date.
An in-flight MOLA diagnostic test (MZ115) was performed on 01-195 (7/14/01) to gain some insight into the MOLA anomaly reported last week. Unfortunately, the MOLA instrument did not assume normal operation, but the data indicate that the MOLA computer is working and that the instrument responds to commanding. We expect further diagnostic testing following the MOLA team's analyses of the most recent data.
All subsystems report good health and status except for the MOLA payload instrument that malfunctioned on 01-181 (6/30/01). The MOLA instrument team is still investigating the anomaly.
There have been 13 uplinks to the spacecraft during the past week, including new star catalogs and ephemeris files, instrument command loads, the MM151 background sequence, and the MZ115 MOLA diagnostic test. 5,500 command files have been radiated to the spacecraft since launch.
A Bistatic Radar experiment is planned for 01-217 (8/5/01). The first of several Delta-DOR experiments will take place on 01-236 (8/24/01).
// end //